The future value of $580 eight years from now at 6 percent. (Round your FV factor to 3 decimal places and final answer to 2 decimal places.)

$______
Answer #1

Given, Present Value = PV = $580

Number of years = n = 8

Rate of return = r = 6% or 0.06

Future Value FV = PV*FV Factor

FV Factor = (1+r)n = (1+0.06)8 = 1.594

=> FV = 580*1.594 = $924.52